OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? negative eight q minus six equals ten negative nine plus s equals negative one four h minus twenty-two equals negative seven seventeen minus f equals 4

OpenStudy (eyust707):

-8q -6 = 10 -9+s = -1 4h - 22 = -7 17-f = 4

OpenStudy (eyust707):

Answering this question requires that you understand how to solve each one for their letter.

OpenStudy (eyust707):

Have you ever been taught how to do that?

OpenStudy (anonymous):

no

OpenStudy (eyust707):

lets look at the first one

OpenStudy (anonymous):

ok

OpenStudy (eyust707):



OpenStudy (eyust707):

first step is to add 6 to both sides